# On the Definition of Intelligence

### v1

@@ 6 Generalization to Unseen Categories
where each set $\{x_{i,1},\,x_{i,2},\,\dots,\,x_{i,m_{i}}\}$ is drawn from a distinct category $C_{i}$, for $i=1,2,\dots,M$. Assume each sample $x_{i,j}$ can be viewed or transformed into a problem–solution pair, depending on the task. Now define
@@
so that each $S_{i}\sim C_{i}$ represents a single sample from a higher-level training space composed of categories. By training on multiple such sets $\{S_{1},S_{2},\dots,S_{M}\}$, the model learns to generalize beyond any single category: it acquires the ability to generate new problem–solution pairs from previously unseen categories. This supports generalizable intelligence: the ability to produce coherent outputs for unseen problems drawn from categories not encountered during training.
A practical example of this can be seen in large language models (LLMs). During training, LLMs observe data drawn from many different implicit categories—spanning languages, domains, and tasks. Although not explicitly trained for each specific case, these models can often generate accurate responses to novel prompts by generalizing across the diverse categorical patterns they have encountered.
Furthermore, suppose a family of admissible distinguishers and an $\varepsilon$-category intelligence metric have been established on a reference category, which we denote as $\mathcal{A}$. If an unseen problem space, formalized as another category $\mathcal{B}$, can be embedded into $\mathcal{A}$ via a functor $F:\mathcal{B}\!\to\!\mathcal{A}$ that is full, faithful, and essentially surjective, then every validated distinguisher in $\mathcal{A}$ can, in principle, be transported along $F$ to $\mathcal{B}$[18], thereby conferring the same theoretical $\varepsilon$-category intelligence bound in the new domain.
@@ 7 Conclusion
This paper proposes a minimal, testable, and general definition of intelligence grounded in the ability to generate samples consistent with a given category. It aims to unify a wide range of intelligent behaviors under a single operational framework, without relying on task-specific or anthropocentric assumptions.
First, in terms of evaluation, by leveraging the principle of sample indistinguishability one can quantify intelligence through the degree of alignment between a system’s outputs and given samples on category attributes. Whether dealing with a generative model, a classifier, or a reinforcement-learning agent, if its outputs are indistinguishable from target-category samples within an acceptable error bound, it can be said to possess the corresponding $\varepsilon$-category intelligence.
@@
Third, the proposed framework naturally supports generalization to novel tasks and unseen categories. Once the system has learned multiple known categories, it can leverage its abstract understanding of category structure to generate consistent samples for new, previously unencountered categories—thereby achieving genuine generalized intelligence. This property aligns with current large-scale pretraining practices.
A promising direction for future research is to elevate this framework using the formal machinery of Category Theory. The “classes” or “concepts” we discuss could be rigorously defined as objects in a category, where the morphisms capture structural similarities. In such a setting, generalization to new problems could be modeled as a functor mapping one problem category to another. This would allow for the formal “transport” of distinguishers and provide a deeper structural account of generalizable intelligence. Exploring this possibility is a key next step for this work.
By focusing on observable generative behavior, this definition lays the foundation for theoretical development, cross-domain evaluation, and the design of safe and general-purpose intelligent systems.
{credits}

### v2

@@ 6 Generalization to Unseen Concepts
where each set $\{x_{i,1},\,x_{i,2},\,\dots,\,x_{i,m_{i}}\}$ is drawn from a distinct concept $k_{i}$ (and its corresponding fibre $C_{k_{i}}$), for $i=1,2,\dots,M$. Assume each entity $x_{i,j}$ can be viewed or transformed into a problem–solution pair, depending on the task. Now define
@@
so that each $S_{i}$ (drawn from $C_{k_{i}}$) represents a single entity from a higher-level training space composed of concepts. By training on multiple such sets $\{S_{1},S_{2},\dots,S_{M}\}$, the model learns to generalize beyond any single concept: it acquires the ability to generate new problem–solution pairs from previously unseen concepts. This supports generalizable intelligence: the ability to produce coherent outputs for unseen problems drawn from concepts not encountered during training.
A practical example of this can be seen in large language models (LLMs). During training, LLMs observe data drawn from many different implicit concepts—spanning languages, domains, and tasks. Although not explicitly trained for each specific case, these models can often generate accurate responses to novel prompts by generalizing across the diverse conceptual patterns they have encountered.
Furthermore, suppose a family of admissible distinguishers and an $\varepsilon$-concept intelligence metric have been established on a reference concept, which we denote as $\mathcal{K}_{A}$. If an unseen problem space, formalized as another concept $\mathcal{K}_{B}$, can be embedded into $\mathcal{K}_{A}$ via a functor $F:\mathcal{K}_{B}\!\to\!\mathcal{K}_{A}$ that is full, faithful, and essentially surjective, then every validated distinguisher in $\mathcal{K}_{A}$ can, in principle, be transported along $F$ to $\mathcal{K}_{B}$[18], thereby conferring the same theoretical $\varepsilon$-concept intelligence bound in the new domain.
@@ 7 Conclusion
This paper proposes a minimal, testable, and general definition of intelligence grounded in the ability to generate entities consistent with a given concept. It aims to unify a wide range of intelligent behaviors under a single operational framework, without relying on task-specific or anthropocentric assumptions.
First, in terms of evaluation, by leveraging the principle of entity indistinguishability one can quantify intelligence through the degree of alignment between a system’s outputs and given entities on concept attributes. Whether dealing with a generative model, a classifier, or a reinforcement-learning agent, if its outputs are indistinguishable from target-concept entities within an acceptable error bound, it can be said to possess the corresponding $\varepsilon$-concept intelligence.
@@
Third, the proposed framework naturally supports generalization to novel tasks and unseen concepts. Once the system has learned multiple known concepts, it can leverage its abstract understanding of concept structure to generate consistent entities for new, previously unencountered concepts—thereby achieving genuine generalized intelligence. This property aligns with current large-scale pretraining practices.
A promising direction for future research is to elevate this framework using the formal machinery of Category Theory. The “concepts” we discuss could be rigorously defined as objects in a category, where the morphisms capture structural similarities. In such a setting, generalization to new problems could be modeled as a functor mapping one problem concept to another. This would allow for the formal “transport” of distinguishers and provide a deeper structural account of generalizable intelligence. Exploring this possibility is a key next step for this work.
By focusing on observable generative behavior, this definition lays the foundation for theoretical development, cross-domain evaluation, and the design of safe and general-purpose intelligent systems. {credits}